OpenStudy (anonymous):

Gabrielle is 12 years older than Mikhail. The sum of their ages is 74 . What is Mikhail's age?

OpenStudy (anonymous):

74 dived by 2 =37-12=25 25

OpenStudy (anonymous):

Mikhail is 31

OpenStudy (anonymous):

Let x = Gabby's age y = MIkhails's age Then \[x + y = 74\] and \[x = 12 + y \] use substitution to find the value of y. once you obtain the value of y, substititute into original equation to find value of x.

OpenStudy (goformit100):

"Welcome to OpenStudy. I can guide regarding this useful site; ask your doubts from me, for it you can message me. Please use the chat for off topic questions. And remember to give the helper a medal, by clicking on "Best Answer". We follow a code of conduct, ( http://openstudy.com/code-of-conduct ). Please take a moment to read it."